OpenStudy (anonymous):

Kelly loves to eat chocolate pizza. Each square inch of chocolate is 30 calories. She orders a slice of chocolate pizza from a pizza with a diameter of 12 inches that is cut with a central angle of 40 degrees. How many calories does she eat?

OpenStudy (anonymous):

A = pi * r^2 r = d/2 r=6 A = pi 36 square inches cross product - 360 degrees = 36pi 40 degrees = x x = 4pi square inches 4pi * 30 = 3 392.93 calories